Johnny Lee Inc. produces a line of small gasoline-powered engines that can be used in a variety of residential machines, ranging from different types of lawnmowers, to snowblowers, to garden tools (such as tillers and weed-whackers). The basic product line consists of three different models, each meant to fill the needs of a different market. Assume you are the cost accountant for this company and that you have been asked by the owner of the company to construct a flexible budget for factory overhead costs, which seem to be growing faster than revenues. Currently, the company uses machine hours (MHs) as the basis for assigning both variable and fixed factory overhead costs to products.

Within the relevant range of output, you determine that the following factory fixed overhead costs per month should occur: engineering support, $16,100; insurance on the manufacturing facility, $6,100; property taxes on the manufacturing facility, $13,100; depreciation on manufacturing equipment, $14,900; and indirect labor costs of supervisory salaries, $15,900, setup labor, $3,500, and materials handling, $3,600. Variable factory overhead costs are budgeted at $76.00 per machine hour, as follows: electricity, $19.00; indirect materials for Material A of $12.00 and for Material B of $15.00; indirect labor—maintenance, $17.00; and production-related supplies, $13.00.

Required:

1. Prepare a flexible budget for Johnny Lee for each of the following monthly levels of machine hours: (a) 5,100, (b) 6,100, and (c) 7,100.

2. Generate an equation to represent, within the relevant range, the factory overhead costs per month for Johnny Lee. Use this equation to estimate monthly total overhead cost for machine hours of 4,100 to 7,100, in increments of 500.
Answer #1

Johnny Lee Inc.

Flexible Budget

Machine Hours 5,100 6,100 7,100
(A) Fixed Overhead
Engineering Support $16,100 $16,100 $16,100
Insurance on Manufacturing Facility $6,100 $6,100 $6,100
Property Taxes on Manufacturing Facility $13,100 $13,100 $13,100
Depreciation $14,900 $14,900 $14,900
Indirect Labor Cost ( Supervisory Salary) $15,900 $15,900 $15,900
Setup Labor $3,500 $3,500 $3,500
Material Handling $3,600 $3,600 $3,600
Total Fixed Overhead $73,200 $73,200 $73,200
(B) Variable Overhead
Electricity @ $19 $96,900 $115,900 $134,900
Indirect Material A @ $12 $61,200 $73,200 $85,200
Indirect Material B @ $15 $76,500 $91,500 $106,500
Indirect Labor (Maintenance) @ $17 $86,700 $103,700 $120,700
Product related supplies @ $13 $66,300 $79,300 $92,300
Total Variable Overhead $387,600 $463,600 $539,600
Total Overhead (A+B) $460,800 $536,800 $612,800

2. Equation

Total Overhead = Fixed Overhead + (Variable Overhead per Machine Hour x Machine Hours)

Total Overhead @ 4,100 Machine Hours = $73,200 + ($76 x 4,100) = $ 73,200 + $ 311.600 = $384,800

Total Overhead @ 4,600 Machine Hours = $73,200 + ($76 x 4,600) = $ 73,200 + $ 349,600 = $422,800

Total Overhead @ 5,100 Machine Hours = $73,200 + ($76 x 5,100) = $ 73,200 + $ 387,600 = $460,800

Total Overhead @ 5,600 Machine Hours = $73,200 + ($76 x 5,600) = $ 73,200 + $ 425,600 = $498,800

Total Overhead @ 6,100 Machine Hours = $73,200 + ($76 x 6,100) = $ 73,200 + $ 463.600 = $536,800

Total Overhead @ 6,600 Machine Hours = $73,200 + ($76 x 6,600) = $ 73,200 + $ 501.600 = $574,800

Total Overhead @ 7,100 Machine Hours = $73,200 + ($76 x 7,100) = $ 73,200 + $ 539.600 = $612,800